OpenStudy (anonymous):

I know that x to the power of 4 times x to the power of 3 equals x to the power of 7 when simplified but what about when adding and subtracting? Would x to power of 2 plus x to the power of 3 be simplified to x to the power of 5?

OpenStudy (anonymous):

x^2 + x^3 cannot be simplified any further the answer to your second question is no.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

You only mess with the exponents when you are multiplying or dividing. When you are adding or subtracting, you do not mess with the exponents and you can only add and subtract number that have the same variable and exponent. When you are adding and subtracting number, they have to be like terms 2^2 + 3^3 cannot be added because they are not like terms because they have different exponents. If they both had the exponent of 2 or 3, we could simplify it more. 2^2 + 2^2 = 4^2 The exponent stays the same

OpenStudy (anonymous):

Thanks very helpful

OpenStudy (phi):

I think Doodle means 2^2 + 2^2 = 2^2 * (1+1)= 2^2 * 2^1= 2^3